through (-5,-7) parallel to 4x+5y=9

Answer by jim_thompson5910(35256)

4x+5y=C

4(-5)+5(-7)=C

-20-35=C

-55=C

C = -55

So the equation is 4x+5y = -55
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------